### v1

# Instructions for *ACL Proceedings

Abstract
Counter-speech (CS) is a key strategy for mitigating online Hate Speech (HS), yet defining the criteria to assess its effectiveness remains an open challenge. We propose a novel computational framework for CS effectiveness classification, grounded in social science concepts. Our framework defines six core dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – which we use to annotate 4,214 CS instances from two benchmark datasets, resulting in a novel linguistic resource released to the community. In addition, we propose two classification strategies, multi-task and dependency-based, achieving strong results (0.94 and 0.96 average F1 respectively on both expert- and user-written CS), outperforming standard baselines, and revealing strong interdependence among dimensions.
Effectiveness of Counter-Speech against Abusive Content: A Multidimensional Annotation and Classification Study
Greta Damo1, Elena Cabrio1 Serena Villata1 1Université Côte d’Azur, CNRS, Inria, I3S, France Correspondence: greta.damo@univ-cotedazur.fr
Content warning: this paper contains unobfuscated examples some readers may find offensive.
1 Introduction
The rise of social media has intensified concerns over online Hate Speech (HS). Counter-speech (CS), non-aggressive responses that challenge hate using facts or alternative views, has emerged as a promising mitigation strategy Schieb and Preuss (2016); Benesch (2014). CS has been studied through various lenses: some see it as direct responses to hate Mathew et al. (2019); Ashida and Komachi (2022), while others argue it should condemn hate and offer victims support and alternative views Vidgen et al. (2021); Hangartner et al. (2021); Vidgen et al. (2020); He et al. (2021). CS is also valued for its potential to shift beliefs and reduce future hate Qian et al. (2019a); Rieger et al. (2018). CS can be written by experts, online users, or chatbots (based on Large Language Models – LLMs). Expert-written CS typically presents structured reasoning and appeals to shared human values, whereas user and LLM responses tend to be more generic and emotional Mun et al. (2023); Tekiroğlu et al. (2022, 2020).
This study evaluates expert- and user-written CS effectiveness using two benchmark datasets Albanyan and Blanco (2022); Chung et al. (2019). Figure 1 shows examples of expert and user CS.
Figure 1: Examples of expert-written and user-written counter-speech.
Current evaluations on CS effectiveness rely on surface-level metrics – coherence, relevance, or non-offensiveness – assuming well-formed CS is inherently effective Chung et al. (2021); Zhu and Bhat (2021); Baheti et al. (2021); Ashida and Komachi (2022); Mun et al. (2023). But concentrating on CS surface form only is not enough to achieve shifting attitudes or reducing harm.
To address this gap, our contribution is twofold: (1) a new multi-dimensional framework for automatic CS effectiveness classification across six dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – from linguistics and argumentation theories, and two classification strategies, multi-task and dependency-based modeling, outperforming standard baselines, achieving average F1 scores of 0.94 and 0.96 respectively, when combining user-written and expert-written CS data. (2) a novel human annotated resource combining two expert- Chung et al. (2019) and user-written Albanyan and Blanco (2022) CS datasets, enriched with the annotation layer with the six dimensions [^1] Datasets, guidelines, and code are available at this link..
2 Related Work
CS, introduced as a mitigation strategy by Susan et al. (2016), has not yet been extensively explored from a detection standpoint. Among the studies, Poudhar et al. (2024) find that a one-shot prompted LLM achieves promising accuracy in classifying manually labeled CS strategies. Mathew et al. (2018) analyze HS/CS tweet pairs, using a boosting algorithm with TF-IDF and lexical features, which achieves 0.77 F1. Mathew et al. (2019) also analyze CS and neutral comments on YouTube, finding that TF-IDF vectors combined with logistic regression achieve 0.73 F1. Ziems et al. (2020) analyze tweets during the COVID-19 crisis, achieving 0.49 F1 for CS. Garland et al. (2020) apply an ensemble learning algorithm to classify HS and CS tweets, achieving F1 scores between 0.76 and 0.97. A key gap in the literature is the lack of standardized evaluation methods for assessing CS. Several studies propose evaluation criteria, but no unified framework emerged. Zheng et al. (2023) assess CS effectiveness through human judgments, focusing on comfort for targets and bystander empathy. Chung et al. (2021) and Zhu and Bhat (2021) similarly rely on human ratings based on qualities such as informativeness, coherence, and stance. Ashida and Komachi (2022) evaluate CS suitability using offensiveness and stance, while Jones et al. (2024) propose an LLM-based evaluation framework, grounded in NGO guidelines, which shows strong alignment with human judgments. Yu et al. (2024) introduce a conversation-level incivility metric, arguing that CS should be judged by its downstream impact on discourse. We advance beyond previous studies by introducing a human-annotated resource that captures fine-grained dimensions of structural and rhetorical message quality, and by developing classification methods that leverage this framework.
3 Effectiveness Metrics
We propose a framework to evaluate CS effectiveness across six dimensions: Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness. Clarity ensures structural coherence and logical flow, allowing to follow texts effectively Persing and Ng (2013); Wachsmuth et al. (2016, 2017); Stapleton and Wu (2015). Evidence strengthens persuasiveness through relevant support such as statistics or expert testimony Wachsmuth et al. (2017); Rinott et al. (2015); Rahimi et al. (2014); Reinard (1988). Emotional Appeal, a rhetorical strategy, shows that empathy and emotion can enhance impact and share Wachsmuth et al. (2017); Aristotle (2007); Govier (2010); Hangartner et al. (2021); Chen et al. (2022); Marone (2015). Rebuttal reflects the ability to address and counter opposing views, improving credibility and persuasion Wachsmuth et al. (2017); Aristotle (2007); Toulmin (1958); Damer (2009); Granger et al. (2009); O’Keefe (1999); Onoda et al. (2015). Audience Adaptation involves tailoring language to the audience’s style and knowledge, improving relatability Pickering and Garrod (2004); Danescu-Niculescu-Mizil et al. (2011); Wang et al. (2014). Fairness promotes respectful tone and avoids dehumanization, supporting ethical discourse Cialdini (1993); Schreier et al. (1995).
Categorical Dimensions (1–3 Likert Scale)
Clarity 3: Clear, logically structured, directly addresses the HS topic.
2: Mostly clear or slightly generic.
1: Vague or incoherent.
Evidence 3: Multiple pieces of information as supporting evidence.
2: One supporting information.
1: No information as evidence.
Rebuttal 3: Multiple rebuttals targeting specific parts of the HS.
2: One rebuttal.
1: No rebuttal.
Fairness 3: Respectful, no swearing, no attacks against the hater.
2: Mostly polite, no attacks, mild offensive.
1: Aggressive, including swearing and personal attacks.
Binary Dimensions (0/1)
Emotional Appeal 1: Language that evokes emotions (either positive or negative).
0: Neutral tone and language.
Audience Adaptation 1: Matches HS tone/complexity.
0: Mismatched HS tone or complexity.
Table 1: Annotation guidelines for CS effectiveness.
Emotional Appeal and Audience Adaptation are binary (0–1) variables, where 1 indicates presence and 0 absence. Similar to prior work simplifying subjective dimensions into binary formats Habernal and Gurevych (2016); Wachsmuth et al. (2017); Oraby et al. (2017); Koszowy et al. (2024), we adopt this approach to reduce ambiguity. Clarity, Evidence, Rebuttal, and Fairness are categorical, using a 1–3 Likert scale, with 3 as the best score. These dimensions rely on more objective indicators (for example, the number of evidence pieces) that capture weak, moderate, and strong instances. Table 1 reports the definitions for the annotation of the effectiveness dimensions.
4 Datasets Description
We extend the annotation on two existing datasets: CONAN Chung et al. (2019) and the Twitter Dataset Albanyan and Blanco (2022). We select CONAN as it is the first expert-curated dataset of HS/CS pairs, widely used as a benchmark. It is a multilingual (English, French, Italian) dataset centered on Islamophobia, containing 4,078 expert-annotated HS/CS pairs. Through translation and paraphrasing, it is expanded to 14,988 pairs. For our experiments, we retain only the English, non-augmented instances, resulting in 3,847 pairs. The Twitter Dataset [^2] Only tweet IDs are publicly available due to privacy policies. We obtained the full data upon request from the authors. contains 5,652 hateful tweets and replies from social media (Twitter/X), capturing the brevity and style typical of online discourse Bonaldi et al. (2024). We focus on the subset labeled as CS, where a clear target is identifiable, yielding 367 HS/CS pairs. Together, these datasets form a combined benchmark of 4,214 HS/CS pairs.
Annotation & IAA. We define annotation guidelines to label CS effectiveness dimensions. To consolidate these guidelines, a pilot study was conducted on 50 HS/CS pairs from each dataset, involving three annotators  [^3] Two female, one other; age group: 21–30; education level: PhD students. with background in computer science. Through different reconciliation phases, the guidelines were improved iteratively. After an initial round, annotators reviewed disagreements to resolve confusion or misinterpretation. Based on their feedback, the guidelines were updated, and a second and final annotation round was conducted. To measure inter-annotator agreement (IAA), Fleiss’s $\kappa$ was used for binary dimensions, Krippendorff’s $\alpha$ for categorical ones, and Percent Agreement for Audience Adaptation due to near-perfect consensus (as other metrics underestimate high non-random agreement). IAA results improved significantly in the second round, after which the final dataset was labeled by one of the annotators. Table 2 shows IAA scores, indicating strong agreement across all dimensions (see Appendix A for pairwise scores). We also show average effectiveness scores, demonstrating that expert-written CS is more effective (avg. 1.62) than user-written CS. CONAN scores higher in all dimensions except emotional appeal, suggesting online users rely more on emotions than evidential reasoning or facts.
Dimension CONAN Twitter
1st 2nd avg. 1st 2nd avg.
emotional 0.30 0.65 0.29 0.40 0.62 0.45
audience 1.00 1.00 0.99 0.99 0.99 1.00
clarity 0.29 0.75 2.72 0.41 0.82 1.96
evidence 0.27 0.73 1.47 0.34 0.75 1.07
rebuttal 0.21 0.78 1.31 0.42 0.62 1.18
fairness 0.63 0.79 2.95 0.53 0.94 2.35
Total AVG 1.62 1.34
Table 2: IAA from two annotation rounds (Fleiss’ $\kappa$ for binary variables, and Krippendorff’s $\alpha$ for categorical ones) and average effectiveness scores.
5 Experimental Setting
We use BERT-based classifiers to predict CS effectiveness scores across the six dimensions, relying on bert-base-uncased Devlin et al. (2018) trained on the combined dataset of 4,214 HS/CS pairs, and cross-dataset evaluations. We use BERT as the sole model to focus on improving it through our framework. By not comparing across different architectures, we isolate the effect of our multi-task and dependency modeling, ensuring that observed gains result from our contributions. We assess the following configurations (Appendix B for details):
Bert_CS: CS text embeddings are used as input. Binary dimensions are trained using Binary Cross-Entropy (BCE) loss function, and categorical ones with Cross-Entropy (CE). Bert_CS_HS: identical to Bert_CS, but the input is the concatenation of HS and CS texts embeddings for more context. Multi-task_divided trains binary dimensions jointly (BCE for Emotional Appeal, Focal Loss for Audience Adaptation) and categorical ones separately (CE), assuming partial correlation. Multi-task_united combines all dimensions in a single model using the sum of their respective loss functions, assuming stronger interdependence. Dependency_matrix_3e introduces a learnable matrix to capture pairwise relationships between dimensions, and is trained for 3 epochs, while Dependency_matrix_6e for 6 epochs to assess the effect of extended training.
6 Results
Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
Train: Combined Dataset
bert_cs 0.67±0.09 0.99±0.00 0.94±0.01 0.96±0.01 0.96±0.01 0.95±0.01 0.91±0.02
bert_cs_hs 0.62±0.12 0.99±0.00 0.94±0.02 0.96±0.01 0.95±0.01 0.94±0.01 0.90±0.02
multi-task_d 0.92±0.02 0.99±0.00 0.92±0.01 0.93±0.01 0.94±0.03 0.92±0.03 0.94±0.01
multi-task_u 0.84±0.04 0.99±0.00 0.93±0.02 0.92±0.02 0.93±0.03 0.93±0.02 0.92±0.01
dep._m._3e 0.80±0.10 0.99±0.00 0.92±0.01 0.90±0.05 0.91±0.03 0.93±0.03 0.91±0.02
dep._m._6e 0.92±0.04 0.99±0.00 0.95±0.01 0.97±0.02 0.96±0.01 0.95±0.02 0.96*±0.01
Train: CONAN $\rightarrow$ Test: Twitter
bert_cs 0.30±0.01 1.00±0.00 0.41±0.04 0.86±0.09 0.77±0.02 0.54±0.21 0.65±0.05
bert_cs_hs 0.30±0.04 1.00±0.00 0.33±0.03 0.84±0.07 0.73±0.05 0.41±0.07 0.60±0.02
multi-task_d 0.49±0.07 1.00±0.00 0.42±0.02 0.91±0.04 0.80±0.02 0.32±0.03 0.65±0.02
multi-task_u 0.52±0.06 1.00±0.00 0.42±0.03 0.93±0.04 0.77±0.02 0.35±0.05 0.66±0.01
dep._m._3e 0.54±0.02 1.00±0.00 0.41±0.02 0.92±0.02 0.77±0.02 0.34±0.05 0.66±0.01
dep._m._6e 0.50±0.05 1.00±0.00 0.42±0.01 0.92±0.03 0.77±0.02 0.38±0.01 0.66*±0.01
Train: Twitter $\rightarrow$ Test: CONAN
bert_cs 0.14±0.00 0.99±0.00 0.74±0.05 0.45±0.06 0.55±0.11 0.92±0.01 0.63±0.03
bert_cs_hs 0.14±0.01 0.99±0.00 0.76±0.06 0.43±0.02 0.55±0.14 0.92±0.02 0.63±0.03
multi-task_d 0.55±0.17 0.99±0.00 0.51±0.24 0.44±0.01 0.60±0.05 0.89±0.13 0.66*±0.04
multi-task_u 0.47±0.23 0.99±0.00 0.38±0.27 0.43±0.02 0.56±0.07 0.72±0.39 0.59±0.09
dep._m._3e 0.40±0.24 0.99±0.00 0.48±0.22 0.43±0.03 0.53±0.09 0.70±0.13 0.59±0.08
dep._m._6e 0.40±0.21 0.99±0.00 0.75±0.04 0.53±0.07 0.47±0.21 0.78±0.14 0.65±0.07
Table 3: F1 scores (± std) for all models with CS embeddings. Statistically significant results are marked with *.
Table 3 shows average F1 scores across five runs (seeds 42, 0, 1, 2, 3) for all model configurations across our CS effectiveness dimensions. We compare BERT baselines, multi-task learning models, and dependency-based architectures (with CS embeddings). Dependency_matrix_6e consistently achieves the highest average F1 score (0.96), outperforming all the others. This performance gain is likely due to the use of a learnable dependency matrix that captures the interrelations between dimensions, allowing the model to exploit cross-dimensional dependencies, unlike models that treat the six dimensions independently. Statistical tests confirm the results differ significantly, especially between the best model, Dependency_m._6e, and all others (Appendix C). Figure 2 visualizes the dependency matrix learned by Dependency_m._6e. Rows represent influenced tasks, columns represent influencing tasks, with higher values indicating stronger influence. Emotional Appeal, for instance, is influenced by Fairness (0.63) and Rebuttal (0.65), while Audience Adaptation is shaped by Clarity (0.98), Rebuttal (0.96), and Fairness (0.84). These inter-dependencies help the model to improve its classification performance across dimensions. Both multi-task models, Multi-task_d. (0.94) and Multi-task_u. (0.92), outperform the BERT baselines, suggesting that the multi-task framework is more effective, with only a slight performance difference between united and divided. Incorporating HS data in Bert_CS_HS (0.90) does not significantly improve performance over Bert_CS (0.91). Emotional Appeal remains the hardest dimension to classify, showing the most variation across models. In this dimension, dependency-based and multi-task models perform best (up to 0.92 for Dependency_m._6e and Multi-task_d.), while BERT baselines lag behind (0.67 and 0.62), suggesting that richer architectures better capture emotional cues. Audience Adaptation scores above 0.99 in all configurations (Appendix E).
Figure 2: Learned Dependency Matrix obtained from dep._m._6e on the combined dataset with seed 42.
Cross-Domain Generalization. In the CONAN $\rightarrow$ Twitter setting, all models show a marked performance drop, with Dependency_m._6e performing best (0.66) alongside Dependency_m._3e and Multi-task_u. Emotional Appeal declines the most, suggesting a domain shift in how emotional content is conveyed. Clarity, Fairness, and Rebuttal scores are also low, reflecting the less structured nature of user-written CS. In the reverse setting, Twitter $\rightarrow$ CONAN, results are similar. Multi-task_d. achieves the highest avg. F1 score (0.66), followed closely by Dependency_m._6e (0.65). The remaining models obtain lower average F1 scores, indicating less robust generalization in this transfer setup. Models trained on expert-written CS seem more adaptable, but Emotional Appeal remains challenging across domains, ranging from 0.14 to 0.55. In contrast, Audience Adaptation remains consistently high. In-Domain Evaluation. Training and testing on the same dataset (Appendix D) confirm that dependency-based models consistently perform best, further supporting the benefit of modeling interrelations between dimensions in classification.
7 Conclusion
As a first contribution of this paper, we release a novel linguistic resource that merges the CONAN and Twitter datasets for CS, manually annotated with six dimensions of CS effectiveness. As a second contribution, we present a framework for automatically classifying CS effectiveness across these dimensions. Results show that the dependency-based model outperforms BERT baselines with 0.96 F1 on both expert- and user-written CS. Multi-task models also surpass BERT with 0.94 F1, confirming that effectiveness dimensions are interdependent and benefit from explicit dependency modeling. Our architecture-agnostic framework improves BERT and generalizes to other transformers. A deep linguistic analysis highlights which dimensions need improvement for CS evaluation.
8 Limitation
Our framework for assessing CS effectiveness dimensions focuses on key aspects such as structural coherence, linguistic strategies, content strength, and rhetorical impact. Given the vast body of literature on counter-speech in social sciences, we focused on the most relevant studies, but acknowledge that further conceptualizations exist.
Another limitation concerns validating the effectiveness of CS with human users. Future work should assess whether CS aligned with our dimensions effectively influences discourse, fosters empathy toward targets, induces belief change in users, and reduces hateful content from the original authors.
Regarding the data, we acknowledge that our dataset is imbalanced, with a higher number of counter-speech examples generated by experts compared to those from online users, a common trait in counter-speech corpora due to the scarcity of high-quality, user-generated responses Susan et al. (2016); Mathew et al. (2019); Chung et al. (2019); Qian et al. (2019b). Additionally, some dimensions, particularly Audience Adaptation, exhibit class imbalance, which may affect model performance. To improve robustness, future studies should leverage larger and more balanced datasets.
Despite these limitations, our study introduces a new evaluation framework for counter-speech effectiveness and provides a model setup that achieves strong classification results. With this work, we aim to support and strengthen the community effort towards the refinement of counter-speech evaluation frameworks and the improvement of automatic approaches for counter-speech effectiveness assessment and generation.
9 Ethical Statement
We are aware that our study involves human exposure to hateful and potentially upsetting content. For this reason, we followed the mitigation strategies proposed by Vidgen et al. (2019) to protect annotators’ mental health. Participants gave their consent to voluntarily participate in the annotation, and were informed that they could withdraw from the study at any time. Moreover, each participant was allocated a small subset of the data (100 examples), and a generous time frame of about three weeks to mitigate prolonged exposure to potentially distressing language. We plan to share the data with the new annotations for research purposes only.
Appendix A Pairwise IAA Scores
Table 4 shows the results of the IAA between each pair of annotators on both CONAN and Twitter datasets, after the first round of annotations. Table 5 shows the pairwise IAA results after the second round. For binary metrics, we use Fleiss’s $\kappa$, while for multi-label metrics we use Krippendorff’s $\alpha$. For the audience adaptation dimension, we use Percent Agreement.
Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
Tw Co Tw Co Tw Co
emotional_appeal 0.35 0.52 0.51 0.24 0.36 0.18
audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
clarity 0.44 0.58 0.56 0.06 0.35 0.30
evidence 0.50 0.70 0.19 0.22 0.21 0.12
rebuttal 0.31 0.32 0.85 0.12 0.28 0.34
fairness 0.56 0.74 0.72 0.57 0.40 0.48
Table 4: First round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
Tw Co Tw Co Tw Co
emotional_appeal 0.88 0.60 0.42 0.85 0.55 0.48
audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
clarity 0.88 0.76 0.86 0.79 0.73 0.71
evidence 0.78 0.78 0.85 0.77 0.64 0.62
rebuttal 1 0.72 0.48 0.90 0.48 0.72
fairness 0.92 0.90 0.95 0.78 0.94 0.70
Table 5: Second round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
Appendix B Additional Experimental Details
In this section, we describe the model configurations used in our experiments. The goal of these experiments is to classify the effectiveness dimension scores for a given counter-speech response. More specifically, we use bert-base-uncased with 110M parameters. In all configurations, the input length is controlled by setting max_length to 128. The batch_size is 16, the random_state is set to 42, 0, 1, 2, and 3, and the learning_rate is 2e-5. All the inputs are tokenized by using BertTokenizer, and the process includes 3 training epochs (6 for two configurations). For performing the experiments, we concatenate both CONAN and Twitter datasets, and then we split it into train, validation, and test subsets, with a percentage of 70-10-20. We also perform cross-validation, by using a single dataset for training and validation, and the other for testing. The experiments were run on one A100 GPU.
We use the following loss functions:
  • Binary Cross-Entropy Loss (BCE): Used for binary classification tasks, it measures the difference between predicted probabilities and actual labels.
  • Cross-Entropy Loss (CE): Generalized for multi-class classification, CE extends BCE to multiple classes by computing the log loss over softmax-normalized probabilities.
  • Focal Loss: an extension of CE/BCE, focal loss addresses class imbalance by down-weighting easy samples and focusing on hard ones. It introduces a modulating factor to reduce the contribution of well-classified examples.
We also provide more details regarding our model contributions:
  • Multi-task_divided. We take the binary dimensions, and we train bert-base-uncased on them separately, then we combine the two losses (BCE and Focal Loss, for imbalanced datasets) to make the predictions. Then, we take the multi-label dimensions, we train them separately with Cross Entropy Loss, and, at the end, we combine all the losses, and we make the predictions on the test set. This configuration assumes that dimensions, for which the losses are added together, are somehow correlated.
  • Multi-task_united. Here, we take all the dimensions together, both binary and multi-label, and we train them with their respective loss functions (BCE, Focal Loss, and Cross-Entropy loss). Then we aggregate all the losses, and we predict the results for all of them. All predictions are done with the same Multi-task bert model. Again, the configuration assumes that dimensions are somehow correlated.
  • Dependency_learnable. All the dimensions are considered together, by training bert-base-uncased on them. Each dimension has its own loss function (BCE for emotional appeal, Focal Loss for audience adaptation, Cross Entropy Loss for all the multi-label dimensions), which are summed up at the end. In addition, there is a new parameter called dependency-matrix. Its rows and columns contain values between 0 and 1, that are weights accounting for the dependency between pairs of dimensions. Its weights are randomly initialized, and their best value is learned during the training phase. We run the model for 3 and 6 epochs.
We experiment with two types of input embeddings: (i) concatenated Hate Speech (HS) and Counter Speech (CS) embeddings to provide broader context and (ii) CS embeddings alone. The latter consistently yields slightly better performance. Consequently, we adopt CS-only embeddings for all configurations in the main part of the paper (Multi-task_d., Multi-task_u., Dependency_m._3e, and Dependency_m._6e). For completeness, Table 6 also reports the average results over five runs for models using concatenated CS+HS embeddings.
Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
Train: Combined Dataset
multitask_d. 0.90±0.02 0.99±0.00 0.92±0.01 0.90±0.01 0.91±0.03 0.91±0.03 0.92±0.01
multitask_u. 0.80±0.04 0.99±0.00 0.92±0.02 0.90±0.02 0.91±0.03 0.92±0.02 0.91±0.01
dep._m._3e 0.78±0.10 0.99±0.00 0.91±0.01 0.88±0.05 0.89±0.03 0.93±0.03 0.90±0.02
dep._m._6e 0.90±0.04 0.99±0.00 0.94±0.01 0.96±0.02 0.96±0.01 0.95±0.02 0.95*±0.01
Train: CONAN $\rightarrow$ Test: Twitter
multitask_d. 0.48±0.07 1.00±0.00 0.38±0.02 0.88±0.04 0.80±0.02 0.33±0.03 0.65±0.02
multitask_u. 0.51±0.06 1.00±0.00 0.40±0.03 0.88±0.04 0.79±0.02 0.35±0.05 0.65±0.01
dep._m._3e 0.56±0.02 1.00±0.00 0.39±0.02 0.90±0.02 0.79±0.02 0.34±0.05 0.66*±0.01
dep._m._6e 0.51±0.05 1.00±0.00 0.38±0.01 0.88±0.03 0.79±0.02 0.34±0.01 0.65±0.01
Train: Twitter $\rightarrow$ Test: CONAN
multitask_d. 0.40±0.17 0.99±0.00 0.49±0.24 0.43±0.01 0.60±0.05 0.86±0.13 0.63±0.04
multitask_u. 0.40±0.23 0.99±0.00 0.45±0.27 0.43±0.02 0.54±0.07 0.70±0.39 0.58±0.09
dep._m._3e 0.32±0.24 0.99±0.00 0.39±0.22 0.43±0.03 0.50±0.09 0.73±0.13 0.56±0.08
dep._m._6e 0.35±0.21 0.99±0.00 0.67±0.04 0.48±0.07 0.43±0.21 0.86±0.14 0.63*±0.07
Table 6: Effectiveness F1 scores (± std) for all models. Statistically significant results are marked with *.
Model Twitter Dataset CONAN Dataset
Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
bert_cs 0.31 $\pm$ 0.04 1.00 $\pm$ 0.00 0.52 $\pm$ 0.11 0.91 $\pm$ 0.03 0.78 $\pm$ 0.06 0.54 $\pm$ 0.09 0.68 $\pm$ 0.02 0.61 $\pm$ 0.15 0.99 $\pm$ 0.00 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.00 0.92 $\pm$ 0.03
bert_cs_hs 0.30 $\pm$ 0.04 1.00 $\pm$ 0.00 0.49 $\pm$ 0.16 0.91 $\pm$ 0.04 0.77 $\pm$ 0.07 0.48 $\pm$ 0.07 0.66 $\pm$ 0.02 0.69 $\pm$ 0.13 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 0.97 $\pm$ 0.02 0.88 $\pm$ 0.19 0.97 $\pm$ 0.03 0.91 $\pm$ 0.05
multi-task_d. 0.51 $\pm$ 0.00 1.00 $\pm$ 0.00 0.45 $\pm$ 0.00 0.97 $\pm$ 0.00 0.62 $\pm$ 0.00 0.49 $\pm$ 0.00 0.67 $\pm$ 0.00 0.95 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00
multi-task_u. 0.47 $\pm$ 0.00 1.00 $\pm$ 0.00 0.31 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.45 $\pm$ 0.00 0.63 $\pm$ 0.00 0.91 $\pm$ 0.00 1.00 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.96 $\pm$ 0.00 0.96 $\pm$ 0.00 0.96 $\pm$ 0.00
dependency_m._3e 0.45 $\pm$ 0.00 1.00 $\pm$ 0.00 0.32 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.51 $\pm$ 0.00 0.64 $\pm$ 0.00 0.86 $\pm$ 0.00 1.00 $\pm$ 0.00 0.97 $\pm$ 0.00 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.95 $\pm$ 0.00
dependency_m._6e 0.59 $\pm$ 0.00 1.00 $\pm$ 0.00 0.41 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.61 $\pm$ 0.00 0.70 $\pm$ 0.00 0.95 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00
Table 7: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS only embeddings. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Model Twitter Dataset CONAN Dataset
Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
bert_cs 0.31 $\pm$ 0.04 1.00 $\pm$ 0.00 0.52 $\pm$ 0.11 0.91 $\pm$ 0.03 0.78 $\pm$ 0.06 0.54 $\pm$ 0.09 0.68 $\pm$ 0.02 0.61 $\pm$ 0.15 0.99 $\pm$ 0.00 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.00 0.92 $\pm$ 0.03
bert_cs_hs 0.30 $\pm$ 0.04 1.00 $\pm$ 0.00 0.49 $\pm$ 0.16 0.91 $\pm$ 0.04 0.77 $\pm$ 0.07 0.48 $\pm$ 0.07 0.66 $\pm$ 0.02 0.69 $\pm$ 0.13 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 0.97 $\pm$ 0.02 0.88 $\pm$ 0.19 0.97 $\pm$ 0.03 0.91 $\pm$ 0.05
multi-task_d. 0.54 $\pm$ 0.19 1.00 $\pm$ 0.00 0.36 $\pm$ 0.24 0.97 $\pm$ 0.07 0.62 $\pm$ 0.25 0.53 $\pm$ 0.15 0.67 $\pm$ 0.05 0.94 $\pm$ 0.03 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.98 $\pm$ 0.03 0.95 $\pm$ 0.01
multi-task_u. 0.47 $\pm$ 0.19 1.00 $\pm$ 0.00 0.31 $\pm$ 0.12 0.97 $\pm$ 0.07 0.56 $\pm$ 0.19 0.36 $\pm$ 0.14 0.61 $\pm$ 0.07 0.89 $\pm$ 0.02 0.99 $\pm$ 0.00 0.91 $\pm$ 0.03 0.94 $\pm$ 0.02 0.96 $\pm$ 0.02 0.97 $\pm$ 0.01 0.92 $\pm$ 0.01
dep._m._3e 0.38 $\pm$ 0.14 1.00 $\pm$ 0.00 0.37 $\pm$ 0.07 0.97 $\pm$ 0.07 0.61 $\pm$ 0.19 0.48 $\pm$ 0.18 0.63 $\pm$ 0.06 0.84 $\pm$ 0.10 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.93 $\pm$ 0.03 0.94 $\pm$ 0.04 0.96 $\pm$ 0.04 0.94 $\pm$ 0.02
dep._m._6e 0.53 $\pm$ 0.16 1.00 $\pm$ 0.00 0.44 $\pm$ 0.20 0.97 $\pm$ 0.07 0.66 $\pm$ 0.21 0.56 $\pm$ 0.15 0.69 $\pm$ 0.03 0.95 $\pm$ 0.03 0.99 $\pm$ 0.00 0.98 $\pm$ 0.02 0.99 $\pm$ 0.01 0.98 $\pm$ 0.02 0.99 $\pm$ 0.02 0.98 $\pm$ 0.01
Table 8: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS+HS embeddings. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Appendix C Statistical Significance Test
The dependency_matrix_6e model consistently achieves the highest average F1 scores across five runs, outperforming all other configurations. A one-way ANOVA test, conducted on the average F1 scores across models, confirms a statistically significant difference in performance (F = 12.92, p < 0.001). To further assess whether dependency_m._6e is significantly better than the rest, we conduct pairwise t-tests between it and each of the other models. Results show that dependency_m._6e significantly outperforms bert_cs (p = 0.000114), bert_cs_hs (p = 0.000010), multi_task_d. (p = 0.003057), multi_task_u. (p = 0.000145), and dependency_m._3e (p = 0.000028) at the 1% significance level. These findings demonstrate that dependency_m._6e provides a statistically significant improvement over other baselines and variants.
Appendix D In-Data Results
Table 7 reports F1 classification results of model configurations, with training and testing on the same dataset for the Twitter Dataset, and for CONAN. The results are the average of the scores obtained from 5 runs with seeds of 42, 0, 1, 2, 3, with CS only embeddings, as the results in the main paper. Table 8 reports results with CS + HS embeddings. In both settings, we confirm the trend presented in the main paper, with the Dependency_matrix_6e achieving the best average F1 score. Furthermore, results show that models trained with only CS embeddings achieve slightly better scores than the ones trained on CS and HS embeddings. This can be due to the fact that adding more context, in this scenario, does not help the model to perform better, but it adds confusion. We also notice that training and testing on expert-written CS (CONAN) seems easier as it achieves higher average F1 scores, compared to training and testing on the Twitter dataset.
Appendix E Error Analysis
In this section, we discuss some examples to better understand the models’ behaviors, in particular on the most difficult class (emotional appeal) and the one achieving almost perfect F1 score (audience adaptation).
Audience Adaptation. It consistently achieves high F1 scores (0.99–1.00) across models, likely due to the similarity in linguistic structure and content between CS and HS. For example, consider the following HS/CS pair from CONAN:
HS: "Islam is not a religion that professes peace and it does not make any good men. It condemns women and encourages delinquency. Lawbreakers are not all of a certain race or religion, but only one race and religion make lawbreakers."
CS: "Today, in many Muslim-majority communities, women receive equal, even higher in some places, treatment as men. For example, many women work as scientists, engineers, lawyers, etc." Both speeches aim to persuade a specific group by matching linguistic complexity, thus ensuring audience adaptation. The CS provides counter-examples to the HS claims, using similarly complex and detailed arguments. This makes it easier for the classifier to recognize this dimension, as both speeches maintain a similar linguistic and argumentative structure. However, this near-perfect performance could reflect data bias, suggesting that audience adaptation might be too easily detected, due to the high class imbalance. Further error analysis should investigate whether models are truly understanding adaptation strategies or simply leveraging surface-level cues.
Emotional Appeal. Emotional appeal is challenging to classify due to its subjective nature and context-dependent expression. For example, consider the following counter-speeches from Twitter and CONAN:
Twitter: "@user this isn’t funny at all. don’t make jokes of the situation when rene herself is affected by it. don’t get her dragged even more and delete this." This CS is direct and uses imperatives to express frustration and urgency, relying on internet culture and social cues like tagging someone and using second-person pronouns ("you").
CONAN: "I thought we’d stopped exporting our convicted criminals last century. Now you advocate exporting people without a trial or a conviction." This example uses sarcasm, rhetorical questioning, and historical references to evoke moral outrage and challenge the opposing argument. The emotional appeal here is more subtle and indirect, contrasting with the explicit frustration in the Twitter example. A model trained solely on Twitter might miss these subtleties found in expert-written CS. To improve emotional appeal classification, domain adaptation or dataset augmentation could help models recognize diverse emotional expressions.

### v2

@@
-Instructions for *ACL Proceedings
+Effectiveness of Counter-Speech against Abusive Content: A Multidimensional Annotation and Classification Study
 Abstract
 Counter-speech (CS) is a key strategy for mitigating online Hate Speech (HS), yet defining the criteria to assess its effectiveness remains an open challenge.
-We propose a novel computational framework for CS effectiveness classification, grounded in social science concepts.
+We propose a novel computational framework for CS effectiveness classification, grounded in linguistics, communication and argumentation concepts.
 Our framework defines six core dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – which we use to annotate 4,214 CS instances from two benchmark datasets, resulting in a novel linguistic resource released to the community.
 In addition, we propose two classification strategies, multi-task and dependency-based, achieving strong results (0.94 and 0.96 average F1 respectively on both expert- and user-written CS), outperforming standard baselines, and revealing strong interdependence among dimensions.
-Effectiveness of Counter-Speech against Abusive Content: A Multidimensional Annotation and Classification Study
-Greta Damo1, Elena Cabrio1 Serena Villata1 1Université Côte d’Azur, CNRS, Inria, I3S, France Correspondence: greta.damo@univ-cotedazur.fr
 Content warning: this paper contains unobfuscated examples some readers may find offensive.
-1 Introduction
+I Introduction
 The rise of social media has intensified concerns over online Hate Speech (HS).
-Counter-speech (CS), non-aggressive responses that challenge hate using facts or alternative views, has emerged as a promising mitigation strategy Schieb and Preuss (2016); Benesch (2014).
-CS has been studied through various lenses: some see it as direct responses to hate Mathew et al.
-(2019); Ashida and Komachi (2022), while others argue it should condemn hate and offer victims support and alternative views Vidgen et al.
-(2021); Hangartner et al.
-(2021); Vidgen et al.
-(2020); He et al.
-(2021).
-CS is also valued for its potential to shift beliefs and reduce future hate Qian et al.
-(2019a); Rieger et al.
-(2018).
+Counter-speech (CS) has emerged as an effective strategy for mitigating harmful online discourse.
+It involves non-aggressive responses, that use credible evidence, factual arguments, and alternative perspectives to challenge hate speech [1, 2].
+Counter-speech has been studied and approached through various lenses: some see it as direct responses to hate speech [3, 4], while others argue it should condemn hate, challenge and offer alternative viewpoints, and offer support to the victim [5, 6, 7, 8].
+CS is also valued for its potential to shift beliefs and reduce future hate [9, 10].
 CS can be written by experts, online users, or chatbots (based on Large Language Models – LLMs).
-Expert-written CS typically presents structured reasoning and appeals to shared human values, whereas user and LLM responses tend to be more generic and emotional Mun et al.
-(2023); Tekiroğlu et al.
-(2022, 2020).
-This study evaluates expert- and user-written CS effectiveness using two benchmark datasets Albanyan and Blanco (2022); Chung et al. (2019). Figure 1 shows examples of expert and user CS.
+Expert-written CS typically presents structured reasoning and appeals to shared human values, whereas user- and LLM-generated responses, in particular, tend to be more generic and emotional, and often lacks depth and relies on surface-level rebuttals [11, 12, 13].
+This study evaluates expert- and user-written CS effectiveness using two benchmark datasets [14, 15].This allows us to compare how the origin of a counter-speech message influences its effectiveness.
+Figure 1 shows an example of a hateful message with user- and expert-written CS responses.
 Figure 1: Examples of expert-written and user-written counter-speech.
-Current evaluations on CS effectiveness rely on surface-level metrics – coherence, relevance, or non-offensiveness – assuming well-formed CS is inherently effective Chung et al.
-(2021); Zhu and Bhat (2021); Baheti et al.
-(2021); Ashida and Komachi (2022); Mun et al.
-(2023).
+Current evaluations of counter-speech effectiveness rely on surface-level metrics – coherence, relevance, or non-offensiveness – assuming well-formed CS is inherently effective [16, 17, 18, 4, 11].
 But concentrating on CS surface form only is not enough to achieve shifting attitudes or reducing harm.
-To address this gap, our contribution is twofold: (1) a new multi-dimensional framework for automatic CS effectiveness classification across six dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – from linguistics and argumentation theories, and two classification strategies, multi-task and dependency-based modeling, outperforming standard baselines, achieving average F1 scores of 0.94 and 0.96 respectively, when combining user-written and expert-written CS data.
-(2) a novel human annotated resource combining two expert- Chung et al.
-(2019) and user-written Albanyan and Blanco (2022) CS datasets, enriched with the annotation layer with the six dimensions [^1] Datasets, guidelines, and code are available at this link..
-2 Related Work
-CS, introduced as a mitigation strategy by Susan et al.
-(2016), has not yet been extensively explored from a detection standpoint.
-Among the studies, Poudhar et al.
-(2024) find that a one-shot prompted LLM achieves promising accuracy in classifying manually labeled CS strategies.
-Mathew et al.
-(2018) analyze HS/CS tweet pairs, using a boosting algorithm with TF-IDF and lexical features, which achieves 0.77 F1.
-Mathew et al.
-(2019) also analyze CS and neutral comments on YouTube, finding that TF-IDF vectors combined with logistic regression achieve 0.73 F1.
-Ziems et al.
-(2020) analyze tweets during the COVID-19 crisis, achieving 0.49 F1 for CS.
-Garland et al.
-(2020) apply an ensemble learning algorithm to classify HS and CS tweets, achieving F1 scores between 0.76 and 0.97.
-A key gap in the literature is the lack of standardized evaluation methods for assessing CS.
+Assessing CS effectiveness is crucial, as it has the potential for significant societal impact, by reducing online hostility and potentially curbing offline violence and discrimination.
+To address this gap, this study introduces a novel structured framework grounded in linguistics, communication, rhetoric, and argumentation theories, helping to evaluate counter-speech effectiveness.
+Our contribution is twofold: (1) a new multi-dimensional framework for automatic counter-speech effectiveness classification across six key dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – from linguistics, communication, rhetoric and argumentation theories, and two classification strategies, multi-task and dependency-based modeling, that outperform standard baselines, achieving average F1 scores of 0.94 and 0.96 respectively, when combining user-written and expert-written CS data.
+(2) a novel human annotated resource comprising two expert- [15] and user-written [14] counter-speech datasets, enriched with the annotation layer with the six dimensions of our effectiveness framework [^1] Datasets, guidelines, and code are available at this link..
+The remainder of the paper is structured as follows.
+Section II reviews prior studies on counter-speech and its evaluation.
+Section III defines the six dimensions used to assess CS effectiveness, tailoring them to existing linguistics and argumentation theories.
+Section IV details the datasets used for the annotation and the classification task, and the annotation process together with the Inter-Annotator-Agreement results.
+Section V outlines the models used, and the experimental setup.
+Section VI presents the results of our classification task, showing comparisons between the baselines and our proposed models.
+Section VII identifies and discusses challenges in specific dimensions and limitations.
+Finally, Section VIII summarizes key findings and future directions.
+II Related Work
+Research on counter-speech has gained increasing attention as a strategy for combating online hate speech.
+However, most studies have focused on hate speech detection and counter-speech generation, while the systematic evaluation of CS effectiveness remains under-explored.
+In this section, we review key contributions on counter-speech analysis, emphasizing the need for a comprehensive CS evaluation framework.
+II-A Hate Speech & Counter Speech Detection
+HS detection is a well-studied NLP task, initially tackled as binary classification on English datasets [19, 20], later extended to fine-grained labels for discrimination and microaggressions [21, 22, 23, 24, 25, 26].
+Early approaches relied on traditional Machine Learning [27, 28], while recent methods leverage transformers like BERT for better contextual modeling [29, 30, 31, 32].
+On the other hand, counter-speech, introduced as a mitigation strategy by [33], has not yet been extensively explored from a detection standpoint.
+Among the studies, [34] propose an early HS/CS detection pipeline, by applying an ensemble learning algorithm to classify HS and CS tweets, achieving F1 scores between 0.76 and 0.97, emphasizing counter-speech complexity due to subjectivity and definitional ambiguity.
+II-B Counter Speech Classification & Evaluation
+Concerning counter-speech classification, [35] find that a one-shot prompted LLM achieves promising accuracy in classifying manually labeled CS strategies.
+[36] analyze HS/CS tweet pairs, using a boosting algorithm with TF-IDF and lexical features, which achieves 0.77 F1.
+[3] also analyze CS and neutral comments on YouTube, finding that TF-IDF vectors combined with logistic regression achieve 0.73 F1.
+[37] conducts a qualitative analysis of counter-speech examples, while [38] analyze tweets during the COVID-19 crisis, achieving an F1 score of 0.49 for CS.
+A key gap in the literature is the lack of standardized evaluation methods for assessing human- or machine-generated counter-speech.
 Several studies propose evaluation criteria, but no unified framework emerged.
-Zheng et al.
-(2023) assess CS effectiveness through human judgments, focusing on comfort for targets and bystander empathy.
-Chung et al.
-(2021) and Zhu and Bhat (2021) similarly rely on human ratings based on qualities such as informativeness, coherence, and stance.
-Ashida and Komachi (2022) evaluate CS suitability using offensiveness and stance, while Jones et al.
-(2024) propose an LLM-based evaluation framework, grounded in NGO guidelines, which shows strong alignment with human judgments.
-Yu et al.
-(2024) introduce a conversation-level incivility metric, arguing that CS should be judged by its downstream impact on discourse.
-We advance beyond previous studies by introducing a human-annotated resource that captures fine-grained dimensions of structural and rhetorical message quality, and by developing classification methods that leverage this framework.
-3 Effectiveness Metrics
-We propose a framework to evaluate CS effectiveness across six dimensions: Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness.
-Clarity ensures structural coherence and logical flow, allowing to follow texts effectively Persing and Ng (2013); Wachsmuth et al.
-(2016, 2017); Stapleton and Wu (2015).
-Evidence strengthens persuasiveness through relevant support such as statistics or expert testimony Wachsmuth et al.
-(2017); Rinott et al.
-(2015); Rahimi et al.
-(2014); Reinard (1988).
-Emotional Appeal, a rhetorical strategy, shows that empathy and emotion can enhance impact and share Wachsmuth et al.
-(2017); Aristotle (2007); Govier (2010); Hangartner et al.
-(2021); Chen et al.
-(2022); Marone (2015).
-Rebuttal reflects the ability to address and counter opposing views, improving credibility and persuasion Wachsmuth et al.
-(2017); Aristotle (2007); Toulmin (1958); Damer (2009); Granger et al.
-(2009); O’Keefe (1999); Onoda et al.
-(2015).
-Audience Adaptation involves tailoring language to the audience’s style and knowledge, improving relatability Pickering and Garrod (2004); Danescu-Niculescu-Mizil et al.
-(2011); Wang et al.
-(2014).
-Fairness promotes respectful tone and avoids dehumanization, supporting ethical discourse Cialdini (1993); Schreier et al.
-(1995).
+[39] assess CS effectiveness through human judgments, focusing on comfort for targets and bystander empathy.
+[16] and [17] similarly rely on human ratings based on qualities such as informativeness, coherence, and stance.
+[4] evaluate CS suitability using offensiveness and stance, while [40] propose an LLM-based evaluation framework, grounded in NGO guidelines, which shows strong alignment with human judgments.
+[41] introduce a conversation-level incivility metric, arguing that CS should be judged by its downstream impact on discourse.
+Finally, [42] focus on evaluating the human-likeness of AI-generated counter-speech.
+While prior work shows growing interest in CS evaluation, we advance beyond previous studies by introducing a human-annotated resource based on our framework that captures fine-grained dimensions of structural coherence, content strength, rhetorical impact, and linguistic strategies, to assess its effectiveness.
+We also develop classification methods that leverage this framework.
+III Effectiveness Metrics
+We propose a framework to evaluate counter-speech effectiveness across six dimensions: Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness.
+These are grounded in communication theory.
+  • Clarity ensures structural coherence and logical flow.
+Effective counter-speech should be clear and logically structured, so that it is easier for the audience to follow the reasoning and understand the main points [43, 44, 45, 46].
+  • Evidence ensures that effective counter-speech is supported by relevant evidence and examples that support the claims and make it more compelling to the audience.
+They should be specific to the topics addressed in the hateful message.
+Studies indicate that presenting multiple pieces of evidence in an argument, such as statistics or expert testimony, enhances persuasiveness and effectiveness [45, 47, 48, 49].
+  • Emotional Appeal is a rhetorical strategy, used to evoke empathy, or other emotions in the audience, which can help strengthen the counter-speech impact.
+The importance of emotions and empathy in designing effective counter-speech has been investigated by numerous studies [45, 50, 51, 6, 52, 53].
+  • Rebuttal shows how using arguments to anticipate and address potential counterarguments, objections, and opposing views improves credibility and strengthens persuasion [45, 50, 54, 55, 56, 57, 58].
+  • Audience Adaptation involves tailoring the language of the counter-speech to a specific audience, taking into account their level of linguistic ability and knowledge, thus improving understanding and relatability.
+From a linguistic point of view, a counter-speech is more effective and understandable if the language employed is similar to the one spoken by the audience [59, 60, 61].
+  • Fairness promotes the use of appropriate and respectful language, respecting freedom of expression, without censoring or dehumanizing the opposing viewpoint.
+Fair language supports ethical discourse and improves effectiveness [62, 63].
+Unlike prior frameworks, focused narrowly on toxicity or engagement, our approach enables fine-grained analysis informed by the theoretical perspectives described in argumentation and communication studies.
+They are divided into binary and categorical metrics.
+TABLE I: Annotation guidelines for CS effectiveness.
 Categorical Dimensions (1–3 Likert Scale)
@@
 0: Mismatched HS tone or complexity.
-Table 1: Annotation guidelines for CS effectiveness.
-Emotional Appeal and Audience Adaptation are binary (0–1) variables, where 1 indicates presence and 0 absence.
-Similar to prior work simplifying subjective dimensions into binary formats Habernal and Gurevych (2016); Wachsmuth et al.
-(2017); Oraby et al.
-(2017); Koszowy et al.
-(2024), we adopt this approach to reduce ambiguity.
-Clarity, Evidence, Rebuttal, and Fairness are categorical, using a 1–3 Likert scale, with 3 as the best score.
-These dimensions rely on more objective indicators (for example, the number of evidence pieces) that capture weak, moderate, and strong instances.
-Table 1 reports the definitions for the annotation of the effectiveness dimensions.
-4 Datasets Description
-We extend the annotation on two existing datasets: CONAN Chung et al.
-(2019) and the Twitter Dataset Albanyan and Blanco (2022).
+Binary metrics.
+Emotional Appeal and Audience Adaptation are binary (0-1), where 1 indicates presence and 0 absence of the dimension.
+Due to their inherent subjectiveness and context-dependency, a binary classification (presence vs.
+absence) seems reasonable.
+Similar to prior work simplifying subjective dimensions into binary formats [64, 45, 65, 66], we adopt this approach to reduce ambiguity and subjectivity.
+Categorical metrics.
+Clarity, Evidence, Rebuttal, and Fairness are categorical dimensions, using a 1–3 Likert scale, with 3 as the best score.
+Compared to the binary variables, these dimensions rely on more objective indicators (for example, the number of pieces of evidence), therefore, a 1-3 scale allows annotators to capture gradual differences, distinguishing between weak, moderate, and strong instances.
+Table I reports a summary of the definitions for the annotation of the effectiveness dimensions.
+IV Datasets Description
+We extend the annotation on two existing datasets: CONAN [15] and the Twitter Dataset [14].
 We select CONAN as it is the first expert-curated dataset of HS/CS pairs, widely used as a benchmark.
@@
 We obtained the full data upon request from the authors.
-contains 5,652 hateful tweets and replies from social media (Twitter/X), capturing the brevity and style typical of online discourse Bonaldi et al.
-(2024).
-We focus on the subset labeled as CS, where a clear target is identifiable, yielding 367 HS/CS pairs.
-Together, these datasets form a combined benchmark of 4,214 HS/CS pairs.
+is a real-world dataset, containing 5,652 hateful tweets and replies obtained from social media (Twitter/X), capturing the brevity and style typical of online discourse.
+We focus on the subset labeled as counter-speech, where a clear target is identifiable, yielding 367 HS/CS pairs.
+Together, these datasets form a combined benchmark of 4,214 HS/CS pairs, used in our experiments.
 Annotation & IAA.
 We define annotation guidelines to label CS effectiveness dimensions.
-To consolidate these guidelines, a pilot study was conducted on 50 HS/CS pairs from each dataset, involving three annotators  [^3] Two female, one other; age group: 21–30; education level: PhD students.
-with background in computer science.
+To consolidate these guidelines [^3] Complete guidelines are available at this link, a pilot study was conducted on 50 HS/CS pairs from each dataset, involving three annotators  [^4] Two female, one other; age group: 21–30; education level: PhD students.
+with background in computer science and linguistics.
 Through different reconciliation phases, the guidelines were improved iteratively.
@@
 To measure inter-annotator agreement (IAA), Fleiss’s $\kappa$ was used for binary dimensions, Krippendorff’s $\alpha$ for categorical ones, and Percent Agreement for Audience Adaptation due to near-perfect consensus (as other metrics underestimate high non-random agreement).
-IAA results improved significantly in the second round, after which the final dataset was labeled by one of the annotators.
-Table 2 shows IAA scores, indicating strong agreement across all dimensions (see Appendix A for pairwise scores).
-We also show average effectiveness scores, demonstrating that expert-written CS is more effective (avg.
-1.62) than user-written CS.
-CONAN scores higher in all dimensions except emotional appeal, suggesting online users rely more on emotions than evidential reasoning or facts.
+Table II shows IAA results among the three annotators.
+The IAA of the second round of annotations improved significantly, compared to the first one, as the annotators had the possibility to discuss and refine their understanding of the guidelines.
+Overall, the final IAA scores indicate strong agreement across all dimensions, so the remaining part of the dataset was labeled by one of the annotators.
+Table II also shows average effectiveness scores across both datasets.
+Expert-written counter-speech obtains an average effectiveness score of 1.62, demonstrating its higher effectiveness compared to user-written counter-speech with an average score of 1.34.
+CONAN scores higher in all dimensions except emotional appeal, suggesting that online users rely more on emotions than evidential reasoning or facts.
+TABLE II: IAA from two annotation rounds (Fleiss’ $\kappa$ for binary variables, and Krippendorff’s $\alpha$ for categorical ones) and average effectiveness scores.
 Dimension CONAN Twitter
@@
 Total AVG 1.62 1.34
-Table 2: IAA from two annotation rounds (Fleiss’ $\kappa$ for binary variables, and Krippendorff’s $\alpha$ for categorical ones) and average effectiveness scores.
-5 Experimental Setting
-We use BERT-based classifiers to predict CS effectiveness scores across the six dimensions, relying on bert-base-uncased Devlin et al.
-(2018) trained on the combined dataset of 4,214 HS/CS pairs, and cross-dataset evaluations.
+V Experimental Setting
+In this section, we describe the model configurations used in our experiments.
+The goal is to predict CS effectiveness scores across six dimensions.
+We use BERT classifiers based on bert-base-uncased [67] to predict the effectiveness scores for a given counter-speech response.
+In all configurations, the max_length is set to 128.
+The batch_size is 16, and the learning_rate is 2e-5.
+We run the experiments using five different seeds (0, 1, 2, 3, 42), and we average the results.
+To perform the experiments, we concatenate both CONAN and Twitter datasets, and then we randomly split it into train, validation, and test subsets, with a percentage of 70-10-20.
+We also perform cross-validation, by using a single dataset for training and validation, and the other for testing.
+The experiments were run on one A100 GPU.
 We use BERT as the sole model to focus on improving it through our framework.
-By not comparing across different architectures, we isolate the effect of our multi-task and dependency modeling, ensuring that observed gains result from our contributions.
-We assess the following configurations (Appendix B for details):
-Bert_CS: CS text embeddings are used as input.
-Binary dimensions are trained using Binary Cross-Entropy (BCE) loss function, and categorical ones with Cross-Entropy (CE).
-Bert_CS_HS: identical to Bert_CS, but the input is the concatenation of HS and CS texts embeddings for more context.
-Multi-task_divided trains binary dimensions jointly (BCE for Emotional Appeal, Focal Loss for Audience Adaptation) and categorical ones separately (CE), assuming partial correlation.
-Multi-task_united combines all dimensions in a single model using the sum of their respective loss functions, assuming stronger interdependence.
-Dependency_matrix_3e introduces a learnable matrix to capture pairwise relationships between dimensions, and is trained for 3 epochs, while Dependency_matrix_6e for 6 epochs to assess the effect of extended training.
-6 Results
-Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
-Train: Combined Dataset
-bert_cs 0.67±0.09 0.99±0.00 0.94±0.01 0.96±0.01 0.96±0.01 0.95±0.01 0.91±0.02
-bert_cs_hs 0.62±0.12 0.99±0.00 0.94±0.02 0.96±0.01 0.95±0.01 0.94±0.01 0.90±0.02
-multi-task_d 0.92±0.02 0.99±0.00 0.92±0.01 0.93±0.01 0.94±0.03 0.92±0.03 0.94±0.01
-multi-task_u 0.84±0.04 0.99±0.00 0.93±0.02 0.92±0.02 0.93±0.03 0.93±0.02 0.92±0.01
-dep._m._3e 0.80±0.10 0.99±0.00 0.92±0.01 0.90±0.05 0.91±0.03 0.93±0.03 0.91±0.02
-dep._m._6e 0.92±0.04 0.99±0.00 0.95±0.01 0.97±0.02 0.96±0.01 0.95±0.02 0.96*±0.01
-Train: CONAN $\rightarrow$ Test: Twitter
-bert_cs 0.30±0.01 1.00±0.00 0.41±0.04 0.86±0.09 0.77±0.02 0.54±0.21 0.65±0.05
-bert_cs_hs 0.30±0.04 1.00±0.00 0.33±0.03 0.84±0.07 0.73±0.05 0.41±0.07 0.60±0.02
-multi-task_d 0.49±0.07 1.00±0.00 0.42±0.02 0.91±0.04 0.80±0.02 0.32±0.03 0.65±0.02
-multi-task_u 0.52±0.06 1.00±0.00 0.42±0.03 0.93±0.04 0.77±0.02 0.35±0.05 0.66±0.01
-dep._m._3e 0.54±0.02 1.00±0.00 0.41±0.02 0.92±0.02 0.77±0.02 0.34±0.05 0.66±0.01
-dep._m._6e 0.50±0.05 1.00±0.00 0.42±0.01 0.92±0.03 0.77±0.02 0.38±0.01 0.66*±0.01
-Train: Twitter $\rightarrow$ Test: CONAN
-bert_cs 0.14±0.00 0.99±0.00 0.74±0.05 0.45±0.06 0.55±0.11 0.92±0.01 0.63±0.03
-bert_cs_hs 0.14±0.01 0.99±0.00 0.76±0.06 0.43±0.02 0.55±0.14 0.92±0.02 0.63±0.03
-multi-task_d 0.55±0.17 0.99±0.00 0.51±0.24 0.44±0.01 0.60±0.05 0.89±0.13 0.66*±0.04
-multi-task_u 0.47±0.23 0.99±0.00 0.38±0.27 0.43±0.02 0.56±0.07 0.72±0.39 0.59±0.09
-dep._m._3e 0.40±0.24 0.99±0.00 0.48±0.22 0.43±0.03 0.53±0.09 0.70±0.13 0.59±0.08
-dep._m._6e 0.40±0.21 0.99±0.00 0.75±0.04 0.53±0.07 0.47±0.21 0.78±0.14 0.65±0.07
-Table 3: F1 scores (± std) for all models with CS embeddings. Statistically significant results are marked with *.
-Table 3 shows average F1 scores across five runs (seeds 42, 0, 1, 2, 3) for all model configurations across our CS effectiveness dimensions.
+By not comparing across different architectures, we isolate the effect of our proposed models, ensuring that observed gains result from our contributions, rather than differences between model architectures.
+We assess the following configurations:
+  • Bert_CS: counter-speech text embeddings are used as input. Binary dimensions are trained using the Binary Cross-Entropy (BCE) loss function, and categorical ones with Cross-Entropy Loss (CE).
+  • Bert_CS_HS: identical to Bert_CS, but the input is the concatenation of hate speech and counter-speech text embeddings, to add more context.
+  • Multi-task_divided.
+Binary dimensions are trained jointly using BCE (for emotional_appeal) and Focal Loss (to mitigate class imbalance for audience_adaptation), while categorical dimensions are trained separately using CE.
+Final predictions are obtained by combining all loss components.
+This setting assumes partial correlation among grouped dimensions.
+  • Multi-task_united.
+All dimensions (binary and categorical) are trained together using their respective loss functions in a single multi-task model.
+The total loss is computed as the sum of individual losses.
+This configuration assumes stronger interdependence among all dimensions.
+  • Dependency_matrix_3e.
+All the dimensions are considered together, with their respective loss function (BCE for emotional_appeal, Focal Loss for audience_adaptation, and Cross Entropy Loss for all the multi-label dimensions), which are summed up at the end.
+Additionally, there is a new parameter, a learnable dependency matrix, to capture pairwise relationships between dimensions.
+Its rows and columns contain values between 0 and 1, that are weights accounting for the dependency between pairs of dimensions.
+Its weights are randomly initialized, and their best value is learned during the training phase for 3 epochs.
+  • Dependency_matrix_6e has the same configuration of Dependency_matrix_3e, but it is trained for 6 epochs to assess the effect of extended training.
+We experiment with two types of input embeddings: (i) concatenated Hate Speech (HS) and Counter Speech (CS) embeddings to provide broader context and (ii) CS embeddings alone.
+The latter consistently yields slightly better performance.
+Consequently, we adopt CS-only embeddings for all our proposed configurations (Multi-task_divided, Multi-task_united, Dependency_matrix_3e, and Dependency_matrix_6e).
+VI Results
+TABLE III: F1 scores (± std) for all models.
+Statistically significant results are marked with *.
+AVG: macro-average.
+Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
+Model Train: Combined Dataset Train: Twitter $\rightarrow$ CONAN Train: CONAN $\rightarrow$ Twitter
+Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table III shows average F1 scores across five runs for all model configurations across our counter-speech effectiveness dimensions.
 We compare BERT baselines, multi-task learning models, and dependency-based architectures (with CS embeddings).
-Dependency_matrix_6e consistently achieves the highest average F1 score (0.96), outperforming all the others.
+The dependency-based model Dependency_matrix_6e consistently achieves the highest F1 score (0.96), outperforming all the others.
 This performance gain is likely due to the use of a learnable dependency matrix that captures the interrelations between dimensions, allowing the model to exploit cross-dimensional dependencies, unlike models that treat the six dimensions independently.
-Statistical tests confirm the results differ significantly, especially between the best model, Dependency_m._6e, and all others (Appendix C).
-Figure 2 visualizes the dependency matrix learned by Dependency_m._6e.
-Rows represent influenced tasks, columns represent influencing tasks, with higher values indicating stronger influence.
-Emotional Appeal, for instance, is influenced by Fairness (0.63) and Rebuttal (0.65), while Audience Adaptation is shaped by Clarity (0.98), Rebuttal (0.96), and Fairness (0.84).
+Statistical tests confirm that the results are significantly different.
+A one-way ANOVA test, conducted on the average F1 scores across models, confirms a statistically significant difference in performance (F = 12.92, p << 0.001), especially between the best model, Dependency_matrix_6e, and all others (pairwise t-tests between it and each of the other models show statistically significant differences at the 1% significance level).
+Figure 2 shows the dependency matrix learned by Dependency_matrix_6e.
+Rows represent influenced tasks, columns represent influencing tasks, with higher values indicating stronger influence (diagonal values are zero as dimensions do not influence themselves).
+We can see that Emotional Appeal is influenced by Fairness (0.63) and Rebuttal (0.65), while Audience Adaptation is shaped by Clarity (0.98), Rebuttal (0.96), and Fairness (0.84).
+Clarity is strongly influenced by Rebuttal (0.95) and Emotional Appeal (0.83); Evidence is influenced by Audience Adaptation (0.98); Fairness is influenced by Emotional Appeal (0.80) and Audience Adaptation (0.95), and Rebuttal by Emotional Appeal (0.7).
 These inter-dependencies help the model to improve its classification performance across dimensions.
-Both multi-task models, Multi-task_d.
-(0.94) and Multi-task_u.
-(0.92), outperform the BERT baselines, suggesting that the multi-task framework is more effective, with only a slight performance difference between united and divided.
-Incorporating HS data in Bert_CS_HS (0.90) does not significantly improve performance over Bert_CS (0.91).
-Emotional Appeal remains the hardest dimension to classify, showing the most variation across models.
-In this dimension, dependency-based and multi-task models perform best (up to 0.92 for Dependency_m._6e and Multi-task_d.), while BERT baselines lag behind (0.67 and 0.62), suggesting that richer architectures better capture emotional cues.
-Audience Adaptation scores above 0.99 in all configurations (Appendix E).
-Figure 2: Learned Dependency Matrix obtained from dep._m._6e on the combined dataset with seed 42.
+Both multi-task models, Multi-task_divided (0.94) and Multi-task_united (0.92), also outperform the BERT baselines, suggesting that the multi-task framework is more effective, with only a slight performance difference between united and divided.
+Fine-tuned BERT models remain competitive: Bert_CS (0.91) slightly outperforms Bert_CS_HS (0.90).
+This suggests that incorporating HS data does not significantly improve classification performance.
+Emotional Appeal remains the hardest dimension to classify, showing the most variation across models, ranging from 0.67 to 0.92 F1 for the combined dataset.
+In this dimension, dependency-based and multi-task models perform best (up to 0.92 for Dependency_matrix_6e and Multi-task_divided), while BERT baselines lag behind (0.67 and 0.62), suggesting that richer architectures better capture emotional cues.
+Audience Adaptation scores above 0.99 in all configurations.
+Although this can be due to class imbalance, we address this concern using focal loss during training, which focuses the training on hard, misclassified examples and reduces the impact of easy, well-classified examples.
+By contrast, Clarity and Evidence show greater variation, pointing to their higher classification difficulty.
+Figure 2: Learned Dependency Matrix obtained from Dependency_matrix_6e on the combined dataset with seed 42.
 Cross-Domain Generalization.
-In the CONAN $\rightarrow$ Twitter setting, all models show a marked performance drop, with Dependency_m._6e performing best (0.66) alongside Dependency_m._3e and Multi-task_u.
+To assess model generalization, we perform cross-domain evaluation.
+In the first setting, models are trained on CONAN (expert-written CS) and tested on Twitter Dataset (user-generated CS).
+All models show a marked performance drop, with Dependency_matrix_6e performing best (0.66) alongside Dependency_matrix_3e and Multi-task_united.
 Emotional Appeal declines the most, suggesting a domain shift in how emotional content is conveyed.
 Clarity, Fairness, and Rebuttal scores are also low, reflecting the less structured nature of user-written CS.
-In the reverse setting, Twitter $\rightarrow$ CONAN, results are similar.
-Multi-task_d.
-achieves the highest avg.
-F1 score (0.66), followed closely by Dependency_m._6e (0.65).
+In the reverse setting, where models are trained on Twitter Dataset (user-written CS) and tested on CONAN (expert-generated CS), results are similar.
+Multi-task_divided achieves the highest F1 score (0.66), followed closely by Dependency_matrix_6e (0.65), which remains highly competitive.
 The remaining models obtain lower average F1 scores, indicating less robust generalization in this transfer setup.
-Models trained on expert-written CS seem more adaptable, but Emotional Appeal remains challenging across domains, ranging from 0.14 to 0.55.
-In contrast, Audience Adaptation remains consistently high.
-In-Domain Evaluation.
-Training and testing on the same dataset (Appendix D) confirm that dependency-based models consistently perform best, further supporting the benefit of modeling interrelations between dimensions in classification.
-7 Conclusion
-As a first contribution of this paper, we release a novel linguistic resource that merges the CONAN and Twitter datasets for CS, manually annotated with six dimensions of CS effectiveness.
-As a second contribution, we present a framework for automatically classifying CS effectiveness across these dimensions.
-Results show that the dependency-based model outperforms BERT baselines with 0.96 F1 on both expert- and user-written CS.
-Multi-task models also surpass BERT with 0.94 F1, confirming that effectiveness dimensions are interdependent and benefit from explicit dependency modeling.
-Our architecture-agnostic framework improves BERT and generalizes to other transformers.
-A deep linguistic analysis highlights which dimensions need improvement for CS evaluation.
-8 Limitation
-Our framework for assessing CS effectiveness dimensions focuses on key aspects such as structural coherence, linguistic strategies, content strength, and rhetorical impact.
-Given the vast body of literature on counter-speech in social sciences, we focused on the most relevant studies, but acknowledge that further conceptualizations exist.
-Another limitation concerns validating the effectiveness of CS with human users.
-Future work should assess whether CS aligned with our dimensions effectively influences discourse, fosters empathy toward targets, induces belief change in users, and reduces hateful content from the original authors.
-Regarding the data, we acknowledge that our dataset is imbalanced, with a higher number of counter-speech examples generated by experts compared to those from online users, a common trait in counter-speech corpora due to the scarcity of high-quality, user-generated responses Susan et al.
-(2016); Mathew et al.
-(2019); Chung et al.
-(2019); Qian et al.
-(2019b).
-Additionally, some dimensions, particularly Audience Adaptation, exhibit class imbalance, which may affect model performance.
-To improve robustness, future studies should leverage larger and more balanced datasets.
-Despite these limitations, our study introduces a new evaluation framework for counter-speech effectiveness and provides a model setup that achieves strong classification results.
-With this work, we aim to support and strengthen the community effort towards the refinement of counter-speech evaluation frameworks and the improvement of automatic approaches for counter-speech effectiveness assessment and generation.
-9 Ethical Statement
-We are aware that our study involves human exposure to hateful and potentially upsetting content.
-For this reason, we followed the mitigation strategies proposed by Vidgen et al.
-(2019) to protect annotators’ mental health.
-Participants gave their consent to voluntarily participate in the annotation, and were informed that they could withdraw from the study at any time.
-Moreover, each participant was allocated a small subset of the data (100 examples), and a generous time frame of about three weeks to mitigate prolonged exposure to potentially distressing language.
-We plan to share the data with the new annotations for research purposes only.
-Appendix A Pairwise IAA Scores
-Table 4 shows the results of the IAA between each pair of annotators on both CONAN and Twitter datasets, after the first round of annotations.
-Table 5 shows the pairwise IAA results after the second round.
-For binary metrics, we use Fleiss’s $\kappa$, while for multi-label metrics we use Krippendorff’s $\alpha$.
-For the audience adaptation dimension, we use Percent Agreement.
-Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
-Tw Co Tw Co Tw Co
-emotional_appeal 0.35 0.52 0.51 0.24 0.36 0.18
-audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
-clarity 0.44 0.58 0.56 0.06 0.35 0.30
-evidence 0.50 0.70 0.19 0.22 0.21 0.12
-rebuttal 0.31 0.32 0.85 0.12 0.28 0.34
-fairness 0.56 0.74 0.72 0.57 0.40 0.48
-Table 4: First round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
-Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
-Tw Co Tw Co Tw Co
-emotional_appeal 0.88 0.60 0.42 0.85 0.55 0.48
-audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
-clarity 0.88 0.76 0.86 0.79 0.73 0.71
-evidence 0.78 0.78 0.85 0.77 0.64 0.62
-rebuttal 1 0.72 0.48 0.90 0.48 0.72
-fairness 0.92 0.90 0.95 0.78 0.94 0.70
-Table 5: Second round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
-Appendix B Additional Experimental Details
-In this section, we describe the model configurations used in our experiments.
-The goal of these experiments is to classify the effectiveness dimension scores for a given counter-speech response.
-More specifically, we use bert-base-uncased with 110M parameters.
-In all configurations, the input length is controlled by setting max_length to 128.
-The batch_size is 16, the random_state is set to 42, 0, 1, 2, and 3, and the learning_rate is 2e-5.
-All the inputs are tokenized by using BertTokenizer, and the process includes 3 training epochs (6 for two configurations).
-For performing the experiments, we concatenate both CONAN and Twitter datasets, and then we split it into train, validation, and test subsets, with a percentage of 70-10-20.
-We also perform cross-validation, by using a single dataset for training and validation, and the other for testing.
-The experiments were run on one A100 GPU.
-We use the following loss functions:
-  • Binary Cross-Entropy Loss (BCE): Used for binary classification tasks, it measures the difference between predicted probabilities and actual labels.
-  • Cross-Entropy Loss (CE): Generalized for multi-class classification, CE extends BCE to multiple classes by computing the log loss over softmax-normalized probabilities.
-  • Focal Loss: an extension of CE/BCE, focal loss addresses class imbalance by down-weighting easy samples and focusing on hard ones.
-It introduces a modulating factor to reduce the contribution of well-classified examples.
-We also provide more details regarding our model contributions:
-  • Multi-task_divided.
-We take the binary dimensions, and we train bert-base-uncased on them separately, then we combine the two losses (BCE and Focal Loss, for imbalanced datasets) to make the predictions.
-Then, we take the multi-label dimensions, we train them separately with Cross Entropy Loss, and, at the end, we combine all the losses, and we make the predictions on the test set.
-This configuration assumes that dimensions, for which the losses are added together, are somehow correlated.
-  • Multi-task_united.
-Here, we take all the dimensions together, both binary and multi-label, and we train them with their respective loss functions (BCE, Focal Loss, and Cross-Entropy loss).
-Then we aggregate all the losses, and we predict the results for all of them.
-All predictions are done with the same Multi-task bert model.
-Again, the configuration assumes that dimensions are somehow correlated.
-  • Dependency_learnable.
-All the dimensions are considered together, by training bert-base-uncased on them.
-Each dimension has its own loss function (BCE for emotional appeal, Focal Loss for audience adaptation, Cross Entropy Loss for all the multi-label dimensions), which are summed up at the end.
-In addition, there is a new parameter called dependency-matrix.
-Its rows and columns contain values between 0 and 1, that are weights accounting for the dependency between pairs of dimensions.
-Its weights are randomly initialized, and their best value is learned during the training phase.
-We run the model for 3 and 6 epochs.
-We experiment with two types of input embeddings: (i) concatenated Hate Speech (HS) and Counter Speech (CS) embeddings to provide broader context and (ii) CS embeddings alone.
-The latter consistently yields slightly better performance.
-Consequently, we adopt CS-only embeddings for all configurations in the main part of the paper (Multi-task_d., Multi-task_u., Dependency_m._3e, and Dependency_m._6e).
-For completeness, Table 6 also reports the average results over five runs for models using concatenated CS+HS embeddings.
-Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
-Train: Combined Dataset
-multitask_d. 0.90±0.02 0.99±0.00 0.92±0.01 0.90±0.01 0.91±0.03 0.91±0.03 0.92±0.01
-multitask_u. 0.80±0.04 0.99±0.00 0.92±0.02 0.90±0.02 0.91±0.03 0.92±0.02 0.91±0.01
-dep._m._3e 0.78±0.10 0.99±0.00 0.91±0.01 0.88±0.05 0.89±0.03 0.93±0.03 0.90±0.02
-dep._m._6e 0.90±0.04 0.99±0.00 0.94±0.01 0.96±0.02 0.96±0.01 0.95±0.02 0.95*±0.01
-Train: CONAN $\rightarrow$ Test: Twitter
-multitask_d. 0.48±0.07 1.00±0.00 0.38±0.02 0.88±0.04 0.80±0.02 0.33±0.03 0.65±0.02
-multitask_u. 0.51±0.06 1.00±0.00 0.40±0.03 0.88±0.04 0.79±0.02 0.35±0.05 0.65±0.01
-dep._m._3e 0.56±0.02 1.00±0.00 0.39±0.02 0.90±0.02 0.79±0.02 0.34±0.05 0.66*±0.01
-dep._m._6e 0.51±0.05 1.00±0.00 0.38±0.01 0.88±0.03 0.79±0.02 0.34±0.01 0.65±0.01
-Train: Twitter $\rightarrow$ Test: CONAN
-multitask_d. 0.40±0.17 0.99±0.00 0.49±0.24 0.43±0.01 0.60±0.05 0.86±0.13 0.63±0.04
-multitask_u. 0.40±0.23 0.99±0.00 0.45±0.27 0.43±0.02 0.54±0.07 0.70±0.39 0.58±0.09
-dep._m._3e 0.32±0.24 0.99±0.00 0.39±0.22 0.43±0.03 0.50±0.09 0.73±0.13 0.56±0.08
-dep._m._6e 0.35±0.21 0.99±0.00 0.67±0.04 0.48±0.07 0.43±0.21 0.86±0.14 0.63*±0.07
-Table 6: Effectiveness F1 scores (± std) for all models. Statistically significant results are marked with *.
+Models trained on expert-written CS (Twitter $\rightarrow$ CONAN) seem more adaptable, but Emotional Appeal remains challenging across domains, ranging from 0.14 to 0.55.
+In contrast, Audience Adaptation remains consistently high, showing robustness across domains, while Fairness remains strong only on expert-written data.
+Overall, our models generalize reasonably well, achieving a good F1 score in these cross-dataset tests.
+This suggests that models learn meaningful, transferable patterns beyond dataset-specific characteristics.
+Given the variability in emotional appeal, domain adaptation strategies or dataset augmentation could help models better capture diverse emotional expressions across different counter-speech sources.
+TABLE IV: F1 scores per effectiveness dimension on Twitter and CONAN datasets.
+AVG: macro-average.
+Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
 Model Twitter Dataset CONAN Dataset
@@
 dependency_m._6e 0.59 $\pm$ 0.00 1.00 $\pm$ 0.00 0.41 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.61 $\pm$ 0.00 0.70 $\pm$ 0.00 0.95 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00
-Table 7: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS only embeddings.
-AVG: macro-average.
-Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
-Model Twitter Dataset CONAN Dataset
-Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
-bert_cs 0.31 $\pm$ 0.04 1.00 $\pm$ 0.00 0.52 $\pm$ 0.11 0.91 $\pm$ 0.03 0.78 $\pm$ 0.06 0.54 $\pm$ 0.09 0.68 $\pm$ 0.02 0.61 $\pm$ 0.15 0.99 $\pm$ 0.00 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.00 0.92 $\pm$ 0.03
-bert_cs_hs 0.30 $\pm$ 0.04 1.00 $\pm$ 0.00 0.49 $\pm$ 0.16 0.91 $\pm$ 0.04 0.77 $\pm$ 0.07 0.48 $\pm$ 0.07 0.66 $\pm$ 0.02 0.69 $\pm$ 0.13 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 0.97 $\pm$ 0.02 0.88 $\pm$ 0.19 0.97 $\pm$ 0.03 0.91 $\pm$ 0.05
-multi-task_d.
-0.54 $\pm$ 0.19 1.00 $\pm$ 0.00 0.36 $\pm$ 0.24 0.97 $\pm$ 0.07 0.62 $\pm$ 0.25 0.53 $\pm$ 0.15 0.67 $\pm$ 0.05 0.94 $\pm$ 0.03 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.98 $\pm$ 0.03 0.95 $\pm$ 0.01
-multi-task_u.
-0.47 $\pm$ 0.19 1.00 $\pm$ 0.00 0.31 $\pm$ 0.12 0.97 $\pm$ 0.07 0.56 $\pm$ 0.19 0.36 $\pm$ 0.14 0.61 $\pm$ 0.07 0.89 $\pm$ 0.02 0.99 $\pm$ 0.00 0.91 $\pm$ 0.03 0.94 $\pm$ 0.02 0.96 $\pm$ 0.02 0.97 $\pm$ 0.01 0.92 $\pm$ 0.01
-dep._m._3e 0.38 $\pm$ 0.14 1.00 $\pm$ 0.00 0.37 $\pm$ 0.07 0.97 $\pm$ 0.07 0.61 $\pm$ 0.19 0.48 $\pm$ 0.18 0.63 $\pm$ 0.06 0.84 $\pm$ 0.10 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.93 $\pm$ 0.03 0.94 $\pm$ 0.04 0.96 $\pm$ 0.04 0.94 $\pm$ 0.02
-dep._m._6e 0.53 $\pm$ 0.16 1.00 $\pm$ 0.00 0.44 $\pm$ 0.20 0.97 $\pm$ 0.07 0.66 $\pm$ 0.21 0.56 $\pm$ 0.15 0.69 $\pm$ 0.03 0.95 $\pm$ 0.03 0.99 $\pm$ 0.00 0.98 $\pm$ 0.02 0.99 $\pm$ 0.01 0.98 $\pm$ 0.02 0.99 $\pm$ 0.02 0.98 $\pm$ 0.01
-Table 8: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS+HS embeddings.
-AVG: macro-average.
-Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
-Appendix C Statistical Significance Test
-The dependency_matrix_6e model consistently achieves the highest average F1 scores across five runs, outperforming all other configurations.
-A one-way ANOVA test, conducted on the average F1 scores across models, confirms a statistically significant difference in performance (F = 12.92, p < 0.001).
-To further assess whether dependency_m._6e is significantly better than the rest, we conduct pairwise t-tests between it and each of the other models.
-Results show that dependency_m._6e significantly outperforms bert_cs (p = 0.000114), bert_cs_hs (p = 0.000010), multi_task_d.
-(p = 0.003057), multi_task_u.
-(p = 0.000145), and dependency_m._3e (p = 0.000028) at the 1% significance level.
-These findings demonstrate that dependency_m._6e provides a statistically significant improvement over other baselines and variants.
-Appendix D In-Data Results
-Table 7 reports F1 classification results of model configurations, with training and testing on the same dataset for the Twitter Dataset, and for CONAN.
-The results are the average of the scores obtained from 5 runs with seeds of 42, 0, 1, 2, 3, with CS only embeddings, as the results in the main paper.
-Table 8 reports results with CS + HS embeddings.
-In both settings, we confirm the trend presented in the main paper, with the Dependency_matrix_6e achieving the best average F1 score.
-Furthermore, results show that models trained with only CS embeddings achieve slightly better scores than the ones trained on CS and HS embeddings.
-This can be due to the fact that adding more context, in this scenario, does not help the model to perform better, but it adds confusion.
-We also notice that training and testing on expert-written CS (CONAN) seems easier as it achieves higher average F1 scores, compared to training and testing on the Twitter dataset.
-Appendix E Error Analysis
-In this section, we discuss some examples to better understand the models’ behaviors, in particular on the most difficult class (emotional appeal) and the one achieving almost perfect F1 score (audience adaptation).
+In-Domain Evaluation.
+Table IV reports F1 classification results of model configurations, with training and testing on the same dataset for the Twitter Dataset and for CONAN.
+Training and testing on the same dataset confirm that dependency-based models consistently perform best across effectiveness dimensions, with Dependency_matrix_6e again achieving the best average F1 score, further supporting the benefit of modeling interrelations between dimensions in classification tasks.
+We also notice that training and testing on expert-written CS (CONAN) seem easier, as it achieves higher average F1 scores, compared to training and testing on the Twitter dataset.
+The results across all three evaluation settings highlight key insights into the classification of counter-speech effectiveness dimensions.
+While emotional appeal presents unique challenges due to its subjectivity, the other dimensions (audience adaptation, clarity, evidence, rebuttal, and fairness) also exhibit distinct patterns that provide valuable directions for model improvement.
+VII Error Analysis and Discussion
+In this section, we discuss some examples to better understand the models’ behaviors on two dimensions, Emotional Appeal (most difficult to classify) and Audience Adaptation (highest F1 scores), and we address some limitations.
 Audience Adaptation.
@@
 For example, consider the following HS/CS pair from CONAN:
-HS: "Islam is not a religion that professes peace and it does not make any good men.
+HS: ”Islam is not a religion that professes peace and it does not make any good men.
 It condemns women and encourages delinquency.
-Lawbreakers are not all of a certain race or religion, but only one race and religion make lawbreakers."
-CS: "Today, in many Muslim-majority communities, women receive equal, even higher in some places, treatment as men.
-For example, many women work as scientists, engineers, lawyers, etc." Both speeches aim to persuade a specific group by matching linguistic complexity, thus ensuring audience adaptation.
+Lawbreakers are not all of a certain race or religion, but only one race and religion make lawbreakers.”
+CS: ”Today, in many Muslim-majority communities, women receive equal, even higher in some places, treatment as men.
+For example, many women work as scientists, engineers, lawyers, etc.” Both speeches aim to persuade a specific group by matching linguistic complexity, thus ensuring audience adaptation.
 The CS provides counter-examples to the HS claims, using similarly complex and detailed arguments.
 This makes it easier for the classifier to recognize this dimension, as both speeches maintain a similar linguistic and argumentative structure.
-However, this near-perfect performance could reflect data bias, suggesting that audience adaptation might be too easily detected, due to the high class imbalance.
+However, this near-perfect performance could reflect data bias, suggesting that audience adaptation might be too easily detected, due to the class imbalance.
 Further error analysis should investigate whether models are truly understanding adaptation strategies or simply leveraging surface-level cues.
@@
 For example, consider the following counter-speeches from Twitter and CONAN:
-Twitter: "@user this isn’t funny at all.
+Twitter: ”@user this isn’t funny at all.
 don’t make jokes of the situation when rene herself is affected by it.
-don’t get her dragged even more and delete this." This CS is direct and uses imperatives to express frustration and urgency, relying on internet culture and social cues like tagging someone and using second-person pronouns ("you").
-CONAN: "I thought we’d stopped exporting our convicted criminals last century.
-Now you advocate exporting people without a trial or a conviction." This example uses sarcasm, rhetorical questioning, and historical references to evoke moral outrage and challenge the opposing argument.
+don’t get her dragged even more and delete this.” This CS is direct and uses imperatives to express frustration and urgency, relying on internet culture and social cues like tagging someone and using second-person pronouns (”you”).
+CONAN: ”I thought we’d stopped exporting our convicted criminals last century.
+Now you advocate exporting people without a trial or a conviction.” This example uses sarcasm, rhetorical questioning, and historical references to evoke moral outrage and challenge the opposing argument.
 The emotional appeal here is more subtle and indirect, contrasting with the explicit frustration in the Twitter example.
@@
 To improve emotional appeal classification, domain adaptation or dataset augmentation could help models recognize diverse emotional expressions.
+Model Limitations.
+We use BERT-based classifiers throughout to ensure that improvements stem from our multi-task and dependency-based framework rather than the underlying model.
+Our goal was not to benchmark architectures but to evaluate the framework itself.
+For example, while Large Language Models (LLMs) are promising, and can achieve good results, they are optimized for generation rather than classification tasks.
+Evaluating LLMs would also require a separate experimental design, which is beyond the scope of this paper, and we leave these comparisons as future work.
+VIII Conclusion
+As a first contribution of this paper, we release a novel linguistic resource to the community, enriching the CONAN and Twitter dataset with manual annotations of six fine-grained dimensions of counter-speech effectiveness: Emotional Appeal, Audience Adaptation, Clarity, Evidence, Rebuttal, and Fairness.
+This new annotation layer provides a richer, more diverse foundation for research on CS evaluation, considering content from both expert- and user-generated sources.
+As a second contribution, we introduce a framework for automatically classifying counter-speech effectiveness across these dimensions.
+Our approach is architecture-agnostic and compatible with a range of transformer backbones.
+Experimental results demonstrate that a dependency-based model, which explicitly models inter-dependencies among the dimensions, significantly outperforms standard BERT baselines, achieving 0.96 F1 score on the combined dataset.
+This includes strong performance on both expert- (CONAN) and user-written (Twitter) CS.
+In addition, our multi-task learning models also surpass BERT with 0.94 F1 score, reinforcing the hypothesis that the six dimensions are interrelated and can be jointly learned more effectively.
+Our cross-domain evaluation reveals substantial domain shifts between CONAN and Twitter, especially in Emotional Appeal, highlighting the challenges of generalization and the need for domain-adaptive modeling techniques.
+Despite this, our framework demonstrates robust improvements over BERT baselines across domains, suggesting its effectiveness can generalize beyond specific model architectures.
+Finally, we conduct a detailed linguistic analysis of model performance across dimensions, revealing which aspects of CS effectiveness remain the most difficult to capture.
+This analysis can inform the design of future CS evaluation systems and guide annotation efforts where additional human oversight is most valuable.
+Acknowledgment
+This work has been partially supported by the French government, through the 3IA Côte d’Azur investments in the project managed by the National Research Agency (ANR) with the reference number ANR-23-IACL-0001.